Could someone show me how to work this problem out step by step so I can understand it. I am totally confused.
A landscaping company placed two orders with a nursery. The first order was for 13 bushes and 4 trees, and totaled $487. The second order was for 6 bushes and 2 trees, and totaled $232. The bill does not list the per-item price. What is the cost of one bush and of one tree?

Thank you so very much!

Answer by Alan3354(69443)

A landscaping company placed two orders with a nursery. The first order was for 13 bushes and 4 trees, and totaled $487. The second order was for 6 bushes and 2 trees, and totaled $232. The bill does not list the per-item price. What is the cost of one bush and of one tree?
---------------
13b + 4t = 487
6b + 2t = 232
--------------
Multiply the 2nd eqn by 2
13b + 4t = 487
12b + 4t = 464
--------------- Subtract
b = 23
Can you finish it?